OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9

OpenStudy (whpalmer4):

What do you get if you add both equations together?

OpenStudy (anonymous):

I did the equation and I think the answer is (1,-3) or infinintly many solutions. :)

OpenStudy (whpalmer4):

What do you get if you add those two equations together?

OpenStudy (whpalmer4):

\[3x-2y=9\]\[-3x+2y=-9\]---------------3x-3x-2y+2y=9-9\]\[0=0\] If you get a true statement like that, it means the lines are identical and there are infinitely many solutions, of which \((1,-3)\) is one. If you get a false statement, such as happens here: \[3x-2y=9\]\[-3x+2y=-10\]-----------------\[3x-3x-2y+2y=9-10\]\[0=-1\]That means the lines are parallel, and there are no solutions.

OpenStudy (anonymous):

so the answer is no solution?

OpenStudy (whpalmer4):

For your problem, we got 0=0 after adding the two equations, so the answer there is infinitely many solutions. I then gave a very similar example where we got 0 = -1, to illustrate a case where the correct answer would be "no solution".